key works for radio wealth, last week the worked as fallows
monday 4 5/8 hrs
tuesday 7 1/6 hrs
wensday 9 3/4 hrs
thursday 8 1/2 hrs
friday 9 1/2 hrs
for how many hours did key worked last week?
show your sulotion Answer by Edwin McCravy(20086) 
key works for radio wealth, last week the worked as fallows
monday 4 5/8 hrs
tuesday 7 1/6 hrs
wensday 9 3/4 hrs
thursday 8 1/2 hrs
friday 9 1/2 hrs
for how many hours did key worked last week?
show your sulotion
The original denominators 8,6,4,2,and 2 will all divide evenly into 24.
So 24 is their least common multiple, or the least common denominator.
4 5/8 = 4 __/24
7 1/6 = 7 __/24
9 3/4 = 9 __/24
8 1/2 = 8 __/24
9 1/2 = 9 __/24
----------
So the rule for fining each new numerator:
1. Divide each original denominator into the new denominator.
2. Take what you get and multiply it by the original numerator.
3. write that for the new numerator:
4 5/8 = 4 __/24 <--- 8 divided into 24 gives 3; 3×5 = 15
7 1/6 = 7 __/24 <--- 6 divided into 24 gives 4; 4×1 = 4
9 3/4 = 9 __/24 <--- 4 divided into 24 gives 6; 6×3 = 18
8 1/2 = 8 __/24 <--- 2 divided into 24 gives 12; 12×1 = 12
9 1/2 = 9 __/24 <--- 2 divided into 24 gives 12; 12×1 = 12
----------
Now we fill in the blanks as above:
4 5/8 = 4 15/24
7 1/6 = 7 4/24
9 3/4 = 9 18/24
8 1/2 = 8 12/24
9 1/2 = 9 12/24
----------
Add up the whole numbers, get 37.
4 5/8 = 4 15/24
7 1/6 = 7 4/24
9 3/4 = 9 18/24
8 1/2 = 8 12/24
9 1/2 = 9 12/24
----------
37
Add up the new numerators, getting 61
and write it over the LCM 24
4 5/8 = 4 15/24
7 1/6 = 7 4/24
9 3/4 = 9 18/24
8 1/2 = 8 12/24
9 1/2 = 9 12/24
----------
37 61/24
Divide 24 into 61
__2_
24)61
48
13
Write the quotient 2 and put the remainder 13 over the divisor 24
That gives 2 13/24
Add the 37:
2 13/24
+37
---------
39 13/24
That's the answer.
Edwin